# Powerformer: A Section-adaptive Transformer for Power Flow Adjustment

### v2

@@ 3.3 The Implementation of Powerformer
$\text{factor}^{t}(\boldsymbol{H}^{(l)},\boldsymbol{Z}_{\Phi}^{(l)},\mathcal{E})=\boldsymbol{V}_{t}^{(l)}\odot\text{softmax}\left(\boldsymbol{A}^{(l)}\right)_{t},$ (13)
$\boldsymbol{A}^{(l)}=[\boldsymbol{a}_{1}^{(l)},...,\boldsymbol{a}_{t}^{(l)},...,\boldsymbol{a}_{k}^{(l)}],\;\,\boldsymbol{a}_{t}^{(l)}=\frac{{\boldsymbol{K}_{t}^{(l)\top}\boldsymbol{Q}^{(l)}}}{\sqrt{d}},$ (14)
$\begin{split}&\boldsymbol{Q}^{(l)}=\mathcal{M}_{Q}^{(l)}(\boldsymbol{Z}_{\Phi}^{(l)}),\\ \boldsymbol{K}_{t}^{(l)}=\mathcal{F}_{K,t}^{(l)}&(\boldsymbol{H}_{t}^{(l)},\mathcal{E}),\;\boldsymbol{V}_{t}^{(l)}=\mathcal{F}_{V,t}^{(l)}(\boldsymbol{H}_{t}^{(l)},\mathcal{E}),\end{split}$ (15)
@@
where $\mathcal{M}_{Q}$, $\mathcal{F}_{K,t}^{(l)}$, $\mathcal{F}_{V,t}^{(l)}$ are the corresponding propagation function for the Query, Key and Value components. Notably, our section-adaptive attention based representation (Eq. 13) diverges from the self-attention based representation (Eq. 3) by employing softmax on the attention matrix (Eq. 14) that considers all factors, rather than the attention matrix of a single head. Furthermore, we employ summation to amalgamate diverse factor representations, rather than concatenation.
Complexity Analysis. The time complexity of Powerformer can be analyzed based on Multi-factor Section-adaptive Attention (MFSA) and Graph Neural Network (GNN) propagation. MFSA is $O(nd)$, where $n$ represents the number of bus nodes and $d$ represents the dimensionality of the embeddings. On the other hand, GNN propagation is $O(nd^{2}+md)$, where $m$ represents the number of transmission lines. As a result, Powerformer exhibits a time complexity of O($nd+nd^{2}+md$).
Joint Training. In order to facilitate the ongoing training of the proposed Powerformer framework, we perform iterative updates on the state representation using a deep reinforcement learning framework. Specifically, in this study, we employ the Dueling DQN method Wang2015DuelingNA to achieve this purpose.
In summary, our Powerformer approach enhances the traditional Transformer’s expressiveness for sectional power system power flow adjustment. This is achieved by incorporating the section information into attention matrix, and utilizing GNN and multi-factor mechanism to enable more effectively system state representations.
4 Case Study
@@
To demonstrate the effectiveness of our proposed Powerformer, comprehensive case studies have been conducted. In this section, we first provide the details of experimental setting. Subsequently, we conduct a comparative analysis of our method with concat and attention-based representation learning methods, as well as several popular adjustment baselines. Finally, we conduct ablation studies to enhance the understanding of our proposed method.
4.1 Experimental Setting
@@
Scenario Generation. To obtain adequate scenarios for training and testing purposes, we first employ two small-scale systems, namely the IEEE 118-bus system and a realistic China 300-bus system. Then, we adopt a very large European 9241-bus system from the PEGASE project into our operations Josz2016ACPF . The detailed scenario generation is summarized as follows:
  • 1.
@@
Table 1: The pre-scheduled lower and upper power bound (MW) of each transmission section in the IEEE 118-bus system, the realistic China 300-bus system, and the European 9241-bus system.
Section $\Phi$ 118-bus System 300-bus System 9241-bus System
Lower Upper Lower Upper Lower Upper
1 90 640 140 1000 15 110
2 50 360 280 1960 840 5880
3 40 290 170 1200 145 1000
4 90 640 240 1680 55 390
5 70 480 460 3200 560 3920
6 45 300 200 1400 360 2520
7 130 880 200 1400 345 2410
8 55 390 480 3360 760 5320
9 130 880 170 1200 130 910
10 90 615 80 590 180 1260
@@
Finally, a total of 1829 scenarios were generated for the IEEE 118-bus system, with 1656 scenarios used for training and 173 for testing. Similarly, the China 300-bus system resulted in 1817 scenarios, including 1637 for training and 180 for testing. The European 9241-bus system produced 2037 scenarios, with 1848 for training and 189 for testing.
Figure 3: The visualization highlights the distribution of Concat, Attention, and Powerformer operations for combining state and transmission section representations on the IEEE 118-bus system.
Refer to caption (a) Concat Refer to caption (b) Attention Refer to caption (c) Powerformer (Ours)
Table 2: The performance of our method and baselines on three different power systems is evaluated based on an average of 5 trials. The performance differences between our method and baselines are indicated in parentheses, with better performance denoted by a higher test success rate and lower test economic cost. Method 118-bus System 300-bus System 9241-bus System Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) Concat 68.44 (-28.17) 626,215  (+4,017) 57.66 (-37.14) 992,048 (+22,366) 37.57  (-5.29) 324,485        (-2) Attention 85.23 (-11.38) 625,128  (+2,930) 85.30  (-9.50) 972,573  (+2,891) 31.13 (-11.73) 325,152    (+665) Powerformer (Ours) 98.19 622,198 97.36 969,682 68.25 324,487
Figure 3: The visualization highlights the distribution of Concat, Attention, and Powerformer operations for combining state and transmission section representations on the IEEE 118-bus system.
Compared Methods and Parameter Settings. We first consider two basic architectures for state representations learning on sectional power system: 1) Concat architecture. We employ a MLP network that takes the concatenated vector of the task encoding vector and state feature vector as input. 2) Attention architecture. We utilize the soft attention network Yang2020MultiTaskRL , which allows agent to dynamically select different network modules based on the section representation. Furthermore, the proposed Powerformer is compared with several popular DRL methods, including Deep Q Network (DQN) MnihKSRVBGRFOPB15 , Double DQN Hasselt2015DeepRL , Dueling DQN Wang2015DuelingNA , Advantage Actor Critic (A2C) Mnih2016AsynchronousMF , Proximal Policy Optimization (PPO) schulman2017proximal .
@@
The proposed Powerformer employs a two-layer networks configuration, specifically the hidden dimension $d$ is 64. Additionally, the section encoder utilizes a two-layer MLP network with a candidate hidden dimension set {128, 64}. To implemente the Powerformer, we utilize the Before Transformer architecture min2022transformer , a type of graph transformer. The Dueling architecture includes a two-layer value network with a dimension of 128 and a one-layer advantage network with a dimension of 128. The section representation is $\boldsymbol{Z}_{\Phi}\in\mathbb{R}^{4m}$, where 4 is the corresponding electrical factors of active power, reactive power, voltage magnitude, and phase angle, $m$ denotes the number of the transmission lines. All network modules in the architecture employ the ReLU activation function. During each training iteration, batches of 64 episodes are sampled from a replay buffer containing 20,000 episodes. The target update interval is set to 100, and the discount factor is 0.9. For exploration, we apply an epsilon-greedy approach, where epsilon linearly decreases from 0.1 to 0.01 over 500,000 training steps and remains constant thereafter.
Evaluations. The evaluation of policies incorporates metrics such as the mean test success rate, mean test economic cost, and mean inference speed across a range of tasks.
4.2 Comparison with Different State Representations
In this subsection, our primary focus revolves around a comprehensive comparison between our novel Powerformer and two commonly used representation learning methods, i.e., concat and attention-based representations. By conducting a comprehensive analysis using the same reinforcement learning framework of Dueling DQN Wang2015DuelingNA , we aim to clearly showcase the remarkable effectiveness and superiority of our proposed Powerformer method.
Figure 4: The learning curves of all methods for 4-section task and 10-section task on three power systems. The experimental results use the median performance and one standard deviation (shaded region) over 5 random seeds to ensure a fair comparison.
Refer to caption (a) Refer to caption (a) IEEE 118-bus System(4-section Task) Refer to caption (b) Realistic 300-bus System(4-section Task) Refer to caption (c) European 9241-bus System(4-section Task) Refer to caption (d) IEEE 118-bus System(10-section Task) Refer to caption (e) Realistic 300-bus System(10-section Task) Refer to caption (f) European 9241-bus System(10-section Task)
Table 3: The performance of our method and baselines in 4-section and 10-section tasks on three different power systems is evaluated based on an average of 5 trials. The performance differences between our method and baselines are indicated in parentheses, with better performance denoted by a higher test success rate and lower test economic cost. Method 118-bus System (4-section Task) 300-bus System (4-section Task) 9241-bus System (4-section Task) Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) DQN 88.97 (-19.66) 693,272  (+1,130) 79.05 (-20.68) 981,695 (+19,605) 52.04 (-16.93) 325,331     (-99) Double DQN 87.63 (-12.04) 693,653  (+1,511) 80.15 (-19.58) 995,038 (+32,948) 55.41 (-13.56) 325,464     (+34) Dueling DQN 91.17  (-8.50) 692,378    (+236) 71.45 (-28.28) 995,279 (+33,189) 52.48 (-16.49) 325,266    (-164) A2C 84.98 (-14.69) 697,853  (+5,711) 93.71  (-6.02) 974,455 (+12,365) 53.07 (-15.90) 325,093    (-337) PPO 26.67 (-73.00) 673,399 (-18,743) 46.88 (-52.85) 937,454 (-24,636) 39.18 (-29.79) 326,614  (+1,184) OPF 83.33 (-16.34) 639,782 (-52,360) 56.78 (-42.59) 956,120  (-5,970) 64.47  (-4.50) 324,995    (-435) Powerformer (Ours) 99.67 692,142 99.73 962,090 68.97 325,430 Method 118-bus System (10-section Task) 300-bus System (10-section Task) 9241-bus System (10-section Task) Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) DQN 61.23 (-35.38) 626,808  (+4,610) 56.74 (-38.06) 991,762 (+22,080) 13.67 (-29.19) 324,658    (+164) Double DQN 53.28 (-43.33) 638,079 (+15,881) 61.24 (-33.49) 990,126 (+20,444) 16.75 (-26.11) 324,792    (+305) Dueling DQN 68.44 (-28.17) 626,215  (+4,017) 57.66 (-37.14) 992,048 (+22,366) 37.57  (-5.29) 324,485        (-2) A2C 80.14 (-16.47) 626,703  (+4,505) 75.61 (-19.19) 976,467  (+6,785) 12.70 (-30.16) 325,000    (+513) PPO 24.28 (-72.33) 596,315 (-25,883) 48.61 (-46.19) 930,581 (-39,101) 19.40 (-23.46) 325,549 (+1,062) OPF 72.25 (-24.36) 606,259 (-15,939) 50.56 (-44.24) 953,349 (-16,333) 51.32  (+8.46) 321,586 (-2,901) Powerformer (Ours) 98.19 622,198 97.36 969,682 68.25 324,487
Figure 4: The learning curves of all methods for 4-section task and 10-section task on three power systems. The experimental results use the median performance and one standard deviation (shaded region) over 5 random seeds to ensure a fair comparison.
The performance results presented in Table 2 showcases a comparison between the Powerformer and two fundamental state encoding methods—Concat and Attention—across three different power system cases. The adopt evaluation is based on two metrics: success rate and economic cost, averaged over five trials. Among these results, it is worth noting that the proposed Powerformer demonstrates the exceptional success rates of 98.19%, 97.36%, and 68.25% respectively, solidifying its position as the standout performer. This is particularly evident in its remarkable performance when applied to large-scale power systems such as the European 9241-bus system. In terms of economic cost, Powerformer tends to be more efficient or on par with the other methods. This suggests that Powerformer can achieve higher performance without a relatively low test economic cost, an important consideration for practical applications.
Furthermore, we utilize the t-SNE toolkit to visually illustrate the distribution of state representations for 10 distinct section tasks on the IEEE 118-bus system in Fig. 3. Our proposed Powerformer approach can effectively cluster 10 different section tasks into 10 distinct central points, while the other two methods cannot accomplish. This also explains the poor performance of the other two state encoding methods, as they couple the transmission section representation and different electrical factors of node features together (Fig. 3a and 3b), making it impossible to execute specific actions.
4.3 Comparison with Different Adjustment Baselines

### v4

@@ 3.3. The Implementation of Powerformer
(13) $\text{factor}^{t}(\bm{H}^{(l)},\bm{Z}_{\Phi}^{(l)},\mathcal{E})=\bm{V}_{t}^{(l)}\odot\text{softmax}\left(\bm{A}^{(l)}\right)_{t},$
(14) $\bm{A}^{(l)}=[\bm{a}_{1}^{(l)},...,\bm{a}_{t}^{(l)},...,\bm{a}_{k}^{(l)}],\;\,\bm{a}_{t}^{(l)}=\frac{{\bm{K}_{t}^{(l)\top}\bm{Q}^{(l)}}}{\sqrt{d}},$
(15) $\begin{split}&\bm{Q}^{(l)}=\mathcal{M}_{Q}^{(l)}(\bm{Z}_{\Phi}^{(l)}),\\ \bm{K}_{t}^{(l)}=\mathcal{F}_{K,t}^{(l)}&(\bm{H}_{t}^{(l)},\mathcal{E}),\;\bm{V}_{t}^{(l)}=\mathcal{F}_{V,t}^{(l)}(\bm{H}_{t}^{(l)},\mathcal{E}),\end{split}$
@@
where $\mathcal{M}_{Q}^{(l)}$, $\mathcal{F}_{K,t}^{(l)}$, $\mathcal{F}_{V,t}^{(l)}$ are the corresponding propagation function for the Query, Key and Value components. Notably, our section-adaptive attention based representation (Eq. 13) diverges from the self-attention based representation (Eq. 3) by employing softmax on the attention matrix (Eq. 14) that considers all factors, rather than the attention matrix of a single head. Furthermore, we employ the summation operation to combine diverse factor representations, rather than concatenation in MHA.
(a) Concat
(b) Attention
(c) Powerformer (Ours)
Figure 3. The visualization highlights the distribution of Concat, Attention, and Powerformer operations for combining state and 10 types of transmission section representations on the IEEE 118-bus system.
Table 1. The performance of our method and baselines on three different power systems is evaluated based on an average of 5 trials. The performance differences between our method and baselines are indicated in parentheses, with better performance denoted by a higher test success rate and lower test economic cost.
Method 118-bus System 300-bus System 9241-bus System
Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$)
Concat 68.44 (-28.17) 626,215  (+4,017) 57.66 (-37.14) 992,048 (+22,366) 37.57  (-5.29) 324,485        (-2)
Attention 85.23 (-11.38) 625,128  (+2,930) 85.30  (-9.50) 972,573  (+2,891) 31.13 (-11.73) 325,152    (+665)
Powerformer (Ours) 98.19 622,198 97.36 969,682 68.25 324,487
Joint Training. To deliver the final dispatch policy based on the proposed Powerformer framework, we adopt the Dueling Deep Q Network (DQN) method [^1] The reason for selecting Dueling DQN is provide in Appendix E to train a deep reinforcement learning agent. Specifically, we introduce a Q-network $Q_{\theta}$ parameterized by $\theta$, which takes the state representation $\text{MFSA}(\mathcal{G})$ as input and estimates the Q-values for each possible dispatch action $d$. The dispatch policy with respect to $Q_{\theta}$ is defined as $\pi_{Q}(\mathcal{G})=\arg\max_{d\in\mathcal{D}}Q_{\theta}(\text{MFSA}(\mathcal{G}),d)$, where $\mathcal{D}$ is the dispatch action space.
During training, the agent first interacts with the environment to generate episodes using the $\epsilon$-greedy policy based on the the current Q-network $Q_{\theta}$. The transition tuples $\{\mathcal{G},d,r,\mathcal{G}^{\prime}\}$ encountered during training are stored in the replay buffer $\mathcal{B}$ (Mnih et al., 2015), where $r$ is the environmental reward. Then the Q-network is trained using mini-batch gradient descent on the temporal difference loss as:
(16) $\mathcal{L}=\mathbb{E}_{\mathcal{B}}\Big{[}Q_{\theta}(\text{MFSA}(\mathcal{G}),d)-(r+\gamma\cdot\max_{d^{\prime}}Q_{\hat{\theta}}(\text{MFSA}(\mathcal{G}^{\prime}),d^{\prime}))\Big{]}^{2},$
where the tuples $\{\mathcal{G},d,r,\mathcal{G}^{\prime}\}$ are sampled from the replay buffer $\mathcal{B}$. DQN uses the target network to make the optimization procedure more stable. The target value $Q_{\hat{\theta}}$ depend on previous network parameters but this dependency is ignored during backpropagation.
3.4. Discussions
Multi-channel graph transformers broadly apply self-attention mechanisms to enhance model performance, our architecture introduces a novel section-adaptive attention mechanism that is adaptive to the fluctuating demands in power system operations. In this subsection, we presents a brief discussion and analysis focusing on three aspects: section-adaptive attention mechanism, state factorization, and GNN integration with Transformers.
Section-adaptive Attention Mechanism. Traditional attention-mask techniques primarily focus on preventing the model from accessing future information during training. However, these techniques typically rely on self-attention, which limit their ability to integrate heterogeneous transmission section information. Our section-adaptive approach differs in that it dynamically optimizes attention distribution based on the current section power flow. This allows for a more sophisticated adaptation to the varying demands of the system, enhancing decision-making for resource allocation.
State Factorization. The state factorization aims to decouple the input into the clear electrical factors (i.e., active power, reactive power, voltage magnitude, and phase angle). They each possess explicit physical meaningful information. This is crucial for understanding the contribution of different factors to the state representation of the power system, which also enhances the interpretability of the achieved representation. As defined in Eq. 11 and Eq. 13, our softmax operates between different factors rather than different nodes. In MHA, the electrical factors are coupling together, making it challenging for each individual attention head to interpret distinct electrical factor significances.
GNN integration with Transformers. (1) For the task of transmission section power system adjustment, it is often to adjust the generators that are close to the section. Therefore, local information is important for the current section adjustment, and we need to use GNN to capture local information. However, as the power flow is adjusted, the overall operating conditions of the power system changes, and it also affects electrical equipment that is far away. Thus, we need to use Transformer to capture global information. (2) By leveraging the capability of GNN, our approach dynamically adapts to topological changes such as maintenance, upgrades, or unforeseen emergencies. For example, when a three-phase short circuit occurs on a transmission line, GNN can accurately capture the changes in local information. This helps to maintain robustness against real-world challenges.
For more detailed discussion, please refer to the Appendix D.
() IEEE 118-bus System (10-section Task)
(a) Realistic 300-bus System (10-section Task)
(b) European 9241-bus System (10-section Task)
Figure 4. The learning curves of all methods for 10-section task on three power systems. The experimental results use the median performance and one standard deviation (shaded region) over 5 random seeds to ensure a fair comparison.
Table 2. The performance of our method and baselines for 10-section tasks on three different power systems is evaluated based on an average of 5 trials. The performance differences between our method and baselines are indicated in parentheses, with better performance denoted by a higher test success rate and lower test economic cost.
Method 118-bus System (10-section Task) 300-bus System (10-section Task) 9241-bus System (10-section Task)
Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$) Success Rate (%) Economic Cost (\$)
DQN 61.23 (-36.96) 626,808  (+4,610) 56.74 (-40.62) 991,762 (+22,080) 13.67 (-54.58) 324,658    (+164)
Double DQN 53.28 (-44.91) 638,079 (+15,881) 61.24 (-36.12) 990,126 (+20,444) 16.75 (-42.14) 324,792    (+305)
Dueling DQN 68.44 (-29.75) 626,215  (+4,017) 57.66 (-39.70) 992,048 (+22,366) 37.57 (-30.68) 324,485        (-2)
A2C 80.14 (-18.05) 626,703  (+4,505) 75.61 (-21.75) 976,467  (+6,785) 12.70 (-55.55) 325,000    (+513)
PPO 24.28 (-73.19) 596,315 (-25,883) 48.61 (-48.75) 930,581 (-39,101) 19.40 (-48.85) 325,549 (+1,062)
[…]
Powerformer (Ours) 98.19 622,198 97.36 969,682 68.25 324,487
Figure 5. The visualization effectively emphasizes the active power of all scenario samples, both before and after adjustment, across all ten transmission sections of the IEEE 118-bus system using our method. Among them, 170 scenarios have been successfully adjusted into the capacity of the transmission section, i.e., 98.19% test success rate reported in Table 1 and 2.
Figure 6. The visualization effectively emphasizes the economic cost of all scenario samples, both before and after adjustment, across all ten transmission sections of the IEEE 118-bus system using our method. The economic costs of almost all samples have decreased, without any instances of cost increase.
4. Experiments
@@
In this section, we first provide the details of experimental setting. Subsequently, we conduct a comparative analysis of our method with commonly used state representation learning methods, as well as several popular DRL-based adjustment baselines (Mnih et al., 2015; Hasselt et al., 2016; Wang et al., 2016; Mnih et al., 2016; Schulman et al., 2017; Dommel and Tinney, 1968) and traditional Optimal Power Flow (OPF) (Dommel and Tinney, 1968). Meanwhile, we use a visualized approach to display the section power flow adjustment, economic cost, and required inference time by using Powerformer adjustments. Finally, we conduct ablation studies to enhance a comprehensive understanding of our proposed method.
4.1. Experimental Setting
@@
Scenario Generation. To obtain adequate scenarios for training and testing purposes, we first employ two small-scale systems, namely the IEEE 118-bus system and a realistic China 300-bus system. Then, we adopt a very large European 9241-bus system from the PEGASE project into our operations (Josz et al., 2016). The detailed scenario generation is provided in Appendix A.
Compared Methods. We first consider two basic architectures of state representations for sectional power system: 1) Concat architecture. We employ a MLP network that takes the concatenated vector of the section encoding vector and state feature vector as input. 2) Attention architecture. We utilize the soft attention network, which allows agent to weight different network modules based on the section representation. Furthermore, the proposed Powerformer is compared with several popular DRL methods, including Deep Q Network (DQN) (Mnih et al., 2015), Double DQN (Hasselt et al., 2016), Dueling DQN (Wang et al., 2016), Advantage Actor Critic (A2C) (Mnih et al., 2016), Proximal Policy Optimization (PPO) (Schulman et al., 2017) and traditional Optimal Power Flow (OPF) (Dommel and Tinney, 1968). Hyperparameter settings are provided in Appendix B.
Evaluations. The evaluation of comparison analysis incorporates metrics such as the test success rate, test economic cost, and inference time. The economic cost is defined as:
(17) $\mathcal{R}(\text{MFSA}(\mathcal{G}),d)=\sum_{i=1}^{N_{G}}(\alpha_{i}(\mathcal{P}_{i}^{G})^{2}+\beta_{i}\mathcal{P}_{i}^{G}+\lambda_{i})$
where $\alpha_{i},\beta_{i}$ and $\lambda_{i}$ are the generation cost coefficients of the generator $i$. $\mathcal{P}_{i}^{G}$ is the active power production of the generator $i$. $N_{G}$ is the number of the generators.
Task Setting. In the context of power flow adjustment across various transmission sections, multi-task learning (Zhang and Yang, 2018; Zhou and Yang, 2023) is suitable to manage power dispatch for simultaneously addresses different scenarios within the power system, where each section’s power flow adjustment treated as a distinct task. In the main manuscript, we establish the 10-section power flow adjustment task on three distinct power systems. More task scenarios are considered in Appendix C.
4.2. Comparison with State Representations
In this subsection, we aim to compare our novel Powerformer with two commonly used representation learning methods (concat and attention mechanism). By conducting a comprehensive analysis using the same reinforcement learning framework of Dueling DQN (Wang et al., 2016), we aim to clearly showcase the remarkable effectiveness and superiority of our proposed Powerformer method.
The performance results presented in Table 1 showcases a comparison between the Powerformer and two fundamental state encoding methods—Concat and Attention—across three different power system cases. The adopt evaluation is based on two metrics: success rate and economic cost, averaged over five trials. Among these results, it is worth noting that the proposed Powerformer demonstrates the exceptional success rates of 98.19%, 97.36%, and 68.25% respectively. This is particularly evident in its remarkable performance when applied to large-scale power systems such as the European 9241-bus system. In terms of economic cost, Powerformer tends to be more efficient or on par with the other methods. This suggests that Powerformer can achieve higher performance without incurring a significant test economic cost, which is a crucial factor to consider in practical applications.
Furthermore, we utilize the t-SNE toolkit to visually illustrate the distribution of state representations for 10 distinct sections on IEEE 118-bus system in Fig. 3. Our proposed Powerformer approach can effectively cluster 10 different section representations into 10 central points, while the other two methods cannot accomplish. This also explains the poor performance of the other two state encoding methods, as they couple the transmission section representation and different electrical factors of node features together (Fig. 3(a) and 3(b)), making the agent impossible to execute specific action.
4.3. Comparison with Adjustment Baselines
To demonstrate the superiority of our method in efficiently and effectively adjusting transmission section power flow, we thoroughly analyze and compare our method with various popular approaches in deep reinforcement learning (DRL) framework, as well as a well-established traditional OPF adjustment method.
As shown in Fig. 4, our proposed method has successfully achieved the best test success rate among all considered DRL architectures on all three power systems. We can analyze and compare the superiority of our method based on the difficulty levels of task, especially in the more challenging large-scale power system case (i.e., Fig. 10(c)), where the effectiveness of our method is highlighted. This is because our method is able to learn better representations of sectional power system states, thus avoiding the coupling of state features. By comparative observing the significant differences on three power systems, we can conclude that when the system case is simple, the task is also relatively simple. As a result, the superiority of our method is not readily apparent.
In Table 2, we provide the comparative results of success rates and economic costs pertaining to power flow adjustment tasks. For the IEEE 118-bus system, Powerformer outperforms all other methods with a success rate of 98.19%, significantly surpassing A2C (80.14%) with a substantial margin. For the China 300-bus system, Powerformer continues to exhibit a satisfying level of performance in comparison to the baseline approaches, accomplishing an impressively high test success rate of 97.36%. For the European-9241 system, our method consistently demonstrates exceptional performance, achieving an impressive success rate of 68.25%; on the other hand, alternative DRL methods fail to surpass the 40% threshold. Additionally, our method consistently demonstrates its exceptional ability to achieve relatively low economic costs across all scenarios.
4.4. Visualization Analysis
To better understand the superiority of the proposed Powerformer, we present illustrations that depict active power and economic cost across the 10 transmission sections of IEEE 118-bus system.
Active Power Comparison. By comparing the active power of all scenario samples before and after Powerformer adjustment across all ten transmission sections on the IEEE 118-bus system, we can find that 170 test scenarios are successfully adjusted into the pre-determined range. Among these cases shown in Fig. 6, the majority involve adjusting the situation where the power flow exceeds the limit to align with the range of the upper bound, thereby guaranteeing the safe operation of the power system and ensuring its long-term stability.
Economic Cost Comparison. By comparing the economic cost of all scenario samples before and after Powerformer adjustment on all ten transmission sections of the IEEE 118-bus system, we can find that the economic cost of all test samples has been reduced. As shown in Fig. 6, after performing further adjustment through Powerformer, the economic cost of nearly every sample has been reduced, with no instances of cost increase. Hence, it can be inferred that our method reliably aims to reduce economic cost.
Figure 7. The inference time comparison of total 173 scenarios on the IEEE 118-bus system. Our method is significantly faster than OPF in all 173 scenarios.
Table 3. The inference time of our method and OPF method on three different power systems. Our method’s inference time is significantly lower than OPF on three datasets.
118-bus System 300-bus System 9241-bus System
OPF 44.822 $\pm$ 38.459 47.668 $\pm$ 28.230 1912.608 $\pm$ 1636.746
Powerformer (Ours) 0.078 $\pm$ 0.151 0.121 $\pm$ 0.176 0.585 $\pm$ 0.602
4.5. Inference Time Analysis
In real-time applications, inference time plays a critical role in large-scale systems, as exemplified by the European 9241-bus system. In such scenarios, power flow adjustment becomes more complex due to the numerous variations that need careful consideration and strategic decision-making. To better understand the superiority of our method in terms of inference speed, we provide Fig. 7 and Table 3 to demonstrate the advantages of our approach when compared to traditional OPF method used in practical scenarios. As shown in Fig. 7, the experimental results demonstrate that, when compared to the traditional OPF method, the time required to successfully achieve the predefined range for each individual test is significantly reduced across a comprehensive set of test scenarios in the IEEE 118-bus system. Moreover, it is notable that our method exhibits a noteworthy reduction in averaged inference time across three different power systems, as evidenced by the compelling numerical statistics presented in Table 3. As a result, our proposed approach exhibits significantly improved time efficiency compared to the traditional OPF method, particularly when dealing with larger system sizes, rendering the OPF method impractical.
4.6. Ablation Study
By conducting thorough ablation studies and performing a detailed analysis, we go deeper insights into our proposed Powerformer (PF) architecture. We carefully assess and quantify the contributions of each component, revealing how they individually impact the overall effectiveness and superiority of the algorithm. Fig. 8 presents the comparison results obtained from an array of ablation experiments conducted on IEEE 118-bus system with 10-section adjustment task. These research findings provide valuable insights, clarifying the roles of each component in improving performance.
Powerformer-S (PF-S). This method only utilizes the representation of Value component $\bm{V}$ (Fig. 2) as the final representation of the system state. In contrast to Powerformer-E, it involves the addition of an additional factorization operation. By analyzing the results in Fig. 8, it is evident that comparing Powerformer-S with Powerformer highlights the inadequacy of relying solely on state representations. This emphasizes the importance of incorporating section information.
Figure 8. The performance of our proposed Powerformer and its ablations on the IEEE 118-bus system.
5. Conclusion
In this paper, we propose Powerformer, a specialized Transformer architecture designed to acquire robust state representations for the DRL agent performing the accurate action. Specifically, our proposed Powerformer incorporates transmission section information into the existing Transformer and develops a section-adaptive attention mechanism. Furthermore, our approach disentangles power system graph data into independent electrical factor graphs to conduct a multi-factor attention mechanism, and employs graph neural networks to extract system state features while considering the graph topological structure of the power system. This allows for the robust representations of the sectional power system state, enabling the development of efficient strategies to effectively manage and adjust transmission section power flow. From the perspective of industrial applications, the utilization of our proposed architecture facilitates controlled power dispatch, leading to improved stability of the power system in real-time power flow adjustment.
Acknowledgements.
This work is supported by the Science and Technology Project of SGCC: Research and Digital Application of High-precision Electric Power Super-scale Pre-trained Visual Model (5108-202218280A-2-395-XG).
Appendix A Scenario Generation
To obtain adequate scenario, we adopt the following steps:
@@ Appendix A Scenario Generation
Finally, a total of 1829 scenarios were generated for the IEEE 118-bus system, with 1656 scenarios used for training and 173 for testing. Similarly, the China 300-bus system resulted in 1817 scenarios, including 1637 for training and 180 for testing. The European 9241-bus system produced 2037 scenarios, with 1848 for training and 189 for testing. The generated scenario cases indicate that the section flow information has exceeded the predefined range. Our goal is to adjust the section flow within the predetermined range through the power generation of generators.
Appendix B parameter setting
@@
The proposed Powerformer employs a two-layer networks configuration, specifically the hidden dimension $d$ is 64. Additionally, the section encoder utilizes a two-layer MLP network with a candidate hidden dimension set {128, 64}. To implemente the Powerformer, we utilize the Before Transformer architecture (Min et al., 2022), a type of graph transformer. The Dueling architecture includes a two-layer value network with a dimension of 128 and a one-layer advantage network with a dimension of 128. The section representation is $\bm{Z}_{\Phi}\in\mathbb{R}^{4m}$, where 4 is the corresponding electrical factors of active power, reactive power, voltage magnitude, and phase angle, $m$ denotes the number of the transmission lines. All network modules in the architecture employ the ReLU activation function. During each training iteration, batches of 64 episodes are sampled from a replay buffer containing 20,000 episodes. The target update interval is set to 100, and the discount factor is 0.9. For exploration, we apply an epsilon-greedy approach, where epsilon linearly decreases from 0.1 to 0.01 over 500,000 training steps and remains constant thereafter.
Appendix C Comparison for both 4-section and 10-section adjustment task